xx Simple Plain Text Output with IOASCII

Simple Plain Text Output with IOASCII

Thomas Radke, Gabrielle Allen

Date

Abstract

Thorn IOASCII provides I/O methods for 1D, 2D, and 3D output of grid arrays and grid functions into files in ASCII format. The precise format is designed for visualisation using the clients xgraphx[1]

or gnuplotx[2].

1 Purpose

Thorn IOASCII registers three I/O methods named IOASCIIx1D, IOASCIIx2D, and IOASCIIx3D with the I/O interface in the flesh.

IOASCIIx1D

creates one-dimensional output of 1D, 2D and 3D grid functions and arrays by slicing through the edge (in the octant case) or center (in all origin centered cases) of the grid in the coordinate directions. In addition, output is provided along a diagonal of the grid, in this case the diagonal always starts at the first grid point (that is, in Fortran notation var(1,1,1)) and the line taken uses grid points increasing by 1 in each direction. [NOTE: The diagonal output is not available for staggered variables]. Output for each direction can be selected individually via parameters. Data is written in ASCII format and goes into files with the names:

<variablexname>x<slice>x[<centerxi>][<centerxj>].{asc|xg}

and for diagonals:

<variablexname>x3Dxdiagonal.{asc|xg}

These files can be processed directly by either xgraph or gnuplot (you can select the style of output via parameter settings).

IOASCIIx2D outputs two-dimensional slices of grid functions and arrays planes. Again, slicing is done through the edge (in the octant case) or center (in all origin centered cases). Data is written in ASCII format and goes into files named

<variablexname>x<plane>x[<center>].{asc|xg}

These files can be visualized by gnuplot using its splot command.

IOASCIIx3D outputs three-dimensional grid functions and arrays as a whole. Data is written in ASCII format and goes into files named

<variablexname>x3D.{asc|xg}

These files can be visualized by gnuplot using its splot command.

2 IOASCII Parameters

General parameters to control all IOASCII’s I/O methods are:

IOASCII::out[123]Dxevery (steerable)

How often to do periodic IOASCII output. If this parameter is set in the parameter file, it will override the setting of the shared IO::outxevery parameter. The output frequency can also be set for individual variables using the outxevery option in an option string appended to the

IOASCII::out[123]Dxvars parameter.

IOASCII::out[123]Dxvars (steerable)

The list of variables to output using an IOASCII I/O method. The variables must be given by their fully qualified variable or group name. The special keyword all requests IOASCII output for all variables. Multiple names must be separated by whitespaces. An option string can be appended in curly braces to the group/variable name. The only option supported so far is outxevery which sets the output frequency for an individual variable (overriding

IOASCII::out[123]Dxevery and IO::outxevery).

IOASCII::out[123]Dxdir

The directory in which to place the IOASCII output files. If the directory doesn’t exist at startup it will be created. If this parameter is set to an empty string IOASCII output will go to the standard output directory as specified in IO::outxdir.

IOASCII::outxformat (steerable)

The output format for floating-point numbers in IOASCII output. This parameter conforms to the format modifier of the C library routine fprintf(3). You can set the format for outputting floating-point numbers (fixed or exponential) as well as their precision (number of digits).

IOASCII::out[123]Dxstyle

The output style for IOASCII output. This parameter chooses between gnuplot- and xgraph-suitable output style, and – for gnuplot – determines whether to also plot the physical time in the output data or not.

Additional parameters to control the IOASCIIx1D I/O method are:

IOASCII::out1Dx[xyzd] (steerable)

Chooses the directions to output in 1D IOASCII format (d stands for diagonal direction).

IOASCII::out1Dx[xyz]linex[xyz], IOASCII::out1Dx[xyz]linex[xyz]i

Chooses the slice centers for 1D lines from IOASCIIx1D. These can be set either in physical or index coordinates. If they are set these parameters will override the default slice center parameters

IO::outx[xyz]line[xyz], IO::outx[xyz]line[xyz]i. Note that the slice center can only be set

for grid functions so far. For CCTKxARRAY variables the slices will always start in the origin, ie. (0, 0, 0) in C ordering.

Additional parameters to control the IOASCIIx2D I/O method are:

IOASCII::out2Dx[xyxzyz]planex[xyz], IOASCII::out2Dx[xyxzyz]planex[xyz]i

Chooses the slice centers for 2D planes from IOASCIIx2D. These can be set either in physical or index coordinates. If they are set these parameters will override the default slice center parameters

IO::outx[xyxzyz]plane[xyz], IO::outx[xyxzyz]plane[xyz]i.

3 Comments

Getting Output from IOBasic’s I/O Mehtods You obtain output by an I/O method by either

setting the appropriate I/O parameters

calling one the routines of the I/O function interface provided by the flesh

For a description of basic I/O parameters and the I/O function interface to invoke I/O methods by application thorns please see the documentation of thorn IOUtil and the flesh.

Building Cactus configurations with IOBasic Since IOASCII uses parameters from IOUtil it also needs this I/O helper thorn be compiled into Cactus and activated at runtime in the ActiveThorns parameter in your parameter file.

4 Examples In this section we include example output for different parameter combinations. Note that all these examples were generated for just a couple of timesteps for an extremely small 3D grid.

4.1 One-dimensional xgraph These options produce data suitable for using with the xgraph visualization client in the format

xxf(t=fixed,x,y=fixed,z=fixed):

xxIOASCII::out1D_everyx=x1
xxIOASCII::out1D_varsx=x"wavetoy::phi"
xxIOASCII::out1D_stylex=x"xgraph"

Output File: phixxx[1][1].xg

xx"Parameterxfilexwavetoy.par
xx"CreatedxSunx19xAugx2001x16:31:43
xx"x-labelxx
xx"y-labelxWAVETOY::phix(yx=x0.0000000000000,xzx=x0.0000000000000),x(yix=x1,xzix=x1)

xx"Timex=x0.0000000000000
xx-0.5000000000000 0.0000000000139
xx0.0000000000000 1.0000000000000
xx0.5000000000000 0.0000000000139

xx"Timex=x0.2500000000000
xx-0.5000000000000 0.0000000000000
xx0.0000000000000 0.4980695458846
xx0.5000000000000 0.0000000000000

xx"Timex=x0.5000000000000
xx-0.5000000000000 0.0019304541362
xx0.0000000000000 -0.7509652270577
xx0.5000000000000 0.0019304541362

4.2 One-dimensional gnuplot

These options produce data suitable for using with the gnuplot visualization client in the format

x f(t,x,y=fixed,z=fixed):

xxIOASCII::out1D_everyx=x1
xxIOASCII::out1D_varsx=x"wavetoy::phi"
xxIOASCII::out1D_stylex=x"gnuplotxf(x)"

xx#Parameterxfilexwavetoy.par
xx#CreatedxSunx19xAugx2001x16:33:07
xx#x-labelxx
xx#y-labelxWAVETOY::phix(yx=x0.0000000000000,xzx=x0.0000000000000),x(yix=x1,xzix=x1)
xx#Timex=x0.0000000000000
xx-0.5000000000000 0.0000000000139
xx0.0000000000000 1.0000000000000
xx0.5000000000000 0.0000000000139
xx#Timex=x0.2500000000000
xx-0.5000000000000 0.0000000000000
xx0.0000000000000 0.4980695458846
xx0.5000000000000 0.0000000000000
xx#Timex=x0.5000000000000
xx-0.5000000000000 0.0019304541362
xx0.0000000000000 -0.7509652270577
xx0.5000000000000 0.0019304541362

4.3 One-dimensional gnuplot (including time)

These options produce data suitable for using with the gnuplot visualization client in the format

t x f(t,x,y=fixed,z=fixed):

xxIOASCII::out1D_everyx=x1
xxIOASCII::out1D_varsx=x"wavetoy::phi"
xxIOASCII::out1D_stylex=x"gnuplotxf(t,x)"

xx#Parameterxfilexwavetoy.par
xx#CreatedxSunx19xAugx2001x16:34:48
xx#x-labelxx
xx#y-labelxWAVETOY::phix(yx=x0.0000000000000,xzx=x0.0000000000000),x(yix=x1,xzix=x1)
xx#Timex=x0.0000000000000
xx0.0000000000000 -0.5000000000000 0.0000000000139
xx0.0000000000000 0.0000000000000 1.0000000000000
xx0.0000000000000 0.5000000000000 0.0000000000139
xx#Timex=x0.2500000000000
xx0.2500000000000 -0.5000000000000 0.0000000000000
xx0.2500000000000 0.0000000000000 0.4980695458846
xx0.2500000000000 0.5000000000000 0.0000000000000
xx#Timex=x0.5000000000000
xx0.5000000000000 -0.5000000000000 0.0019304541362
xx0.5000000000000 0.0000000000000 -0.7509652270577
xx0.5000000000000 0.5000000000000 0.0019304541362

4.4 Two-dimensional gnuplot

These options produce data suitable for using with the gnuplot visualization client in the format

x y f(t,x,y,z=fixed):

xxIOASCII::out2D_everyx=x1
xxIOASCII::out2D_varsx=x"wavetoy::phi"
xxIOASCII::out2D_stylex=x"gnuplotxf(x,y)"

xx#Parameterxfilexwavetoy.par
xx#CreatedxSunx19xAugx2001x16:31:43
xx#x-labelxx
xx#y-labelxy
xx#z-labelxWAVETOY::phix(zx=x0.0000000000000),x(zix=x1)

xx#Timex=x0.0000000000000
xx-0.5000000000000 -0.5000000000000 0.0000000000000
xx0.0000000000000 -0.5000000000000 0.0000000000139
xx0.5000000000000 -0.5000000000000 0.0000000000000
xx-0.5000000000000 0.0000000000000 0.0000000000139
xx0.0000000000000 0.0000000000000 1.0000000000000
xx0.5000000000000 0.0000000000000 0.0000000000139
xx-0.5000000000000 0.5000000000000 0.0000000000000
xx0.0000000000000 0.5000000000000 0.0000000000139
xx0.5000000000000 0.5000000000000 0.0000000000000

xx#Timex=x0.2500000000000
xx-0.5000000000000 -0.5000000000000 0.0000000000000
xx0.0000000000000 -0.5000000000000 0.0000000000000
xx0.5000000000000 -0.5000000000000 0.0000000000000
xx-0.5000000000000 0.0000000000000 0.0000000000000
xx0.0000000000000 0.0000000000000 0.4980695458846
xx0.5000000000000 0.0000000000000 0.0000000000000
xx-0.5000000000000 0.5000000000000 0.0000000000000
xx0.0000000000000 0.5000000000000 0.0000000000000
xx0.5000000000000 0.5000000000000 0.0000000000000

xx#Timex=x0.5000000000000
xx-0.5000000000000 -0.5000000000000 0.0000000008425
xx0.0000000000000 -0.5000000000000 0.0019304541362
xx0.5000000000000 -0.5000000000000 0.0000000008425
xx-0.5000000000000 0.0000000000000 0.0019304541362
xx0.0000000000000 0.0000000000000 -0.7509652270577
xx0.5000000000000 0.0000000000000 0.0019304541362

xx-0.5000000000000 0.5000000000000 0.0000000008425
xx0.0000000000000 0.5000000000000 0.0019304541362
xx0.5000000000000 0.5000000000000 0.0000000008425

4.5 Two-dimensional gnuplot (including time)

These options produce data suitable for using with the gnuplot visualization client in the format

t x y f(t,x,y,z=fixed):

xxIOASCII::out2D_everyx=x1
xxIOASCII::out2D_varsx=x"wavetoy::phi"
xxIOASCII::out2D_stylex=x"gnuplotxf(t,x,y)"

xx#Parameterxfilexwavetoy.par
xx#CreatedxSunx19xAugx2001x16:33:07
xx#x-labelxx
xx#y-labelxy
xx#z-labelxWAVETOY::phix(zx=x0.0000000000000),x(zix=x1)

xx#Timex=x0.0000000000000
xx0.0000000000000 -0.5000000000000 -0.5000000000000 0.0000000000000
xx0.0000000000000 0.0000000000000 -0.5000000000000 0.0000000000139
xx0.0000000000000 0.5000000000000 -0.5000000000000 0.0000000000000
xx0.0000000000000 -0.5000000000000 0.0000000000000 0.0000000000139
xx0.0000000000000 0.0000000000000 0.0000000000000 1.0000000000000
xx0.0000000000000 0.5000000000000 0.0000000000000 0.0000000000139
xx0.0000000000000 -0.5000000000000 0.5000000000000 0.0000000000000
xx0.0000000000000 0.0000000000000 0.5000000000000 0.0000000000139
xx0.0000000000000 0.5000000000000 0.5000000000000 0.0000000000000

xx#Timex=x0.2500000000000
xx0.2500000000000 -0.5000000000000 -0.5000000000000 0.0000000000000
xx0.2500000000000 0.0000000000000 -0.5000000000000 0.0000000000000
xx0.2500000000000 0.5000000000000 -0.5000000000000 0.0000000000000
xx0.2500000000000 -0.5000000000000 0.0000000000000 0.0000000000000
xx0.2500000000000 0.0000000000000 0.0000000000000 0.4980695458846
xx0.2500000000000 0.5000000000000 0.0000000000000 0.0000000000000
xx0.2500000000000 -0.5000000000000 0.5000000000000 0.0000000000000
xx0.2500000000000 0.0000000000000 0.5000000000000 0.0000000000000
xx0.2500000000000 0.5000000000000 0.5000000000000 0.0000000000000

xx#Timex=x0.5000000000000
xx0.5000000000000 -0.5000000000000 -0.5000000000000 0.0000000008425
xx0.5000000000000 0.0000000000000 -0.5000000000000 0.0019304541362
xx0.5000000000000 0.5000000000000 -0.5000000000000 0.0000000008425
xx0.5000000000000 -0.5000000000000 0.0000000000000 0.0019304541362
xx0.5000000000000 0.0000000000000 0.0000000000000 -0.7509652270577
xx0.5000000000000 0.5000000000000 0.0000000000000 0.0019304541362

xx0.5000000000000 -0.5000000000000 0.5000000000000 0.0000000008425
xx0.5000000000000 0.0000000000000 0.5000000000000 0.0019304541362
xx0.5000000000000 0.5000000000000 0.5000000000000 0.0000000008425

References

 [1]xxxhttp://www.cactuscode.org/VizTools/xgraph.html,
   http://jean-luc.aei.mpg.de/Codes/xgraph/

 [2]xxxhttp://www.cactuscode.org/VizTools/Gnuplot.html, http://www.gnuplot.info

5 Parameters

11out1dxd 10Scope: privatex01 BOOLEAN

11Description: Do 1D IOASCII output in the diagonal-direction

11                   10  x01  Default: yes

11

11out1dxdir  10Scope: privatex01  STRING

11Description: Output directory for 1D IOASCII files, overrides IO::outxdir

11 Range 10 x01 Default: (none)
11.+  10A valid directory name
11ˆ$  10An empty string to choose the default from IO::outxdir

11

11out1dxevery  10Scope: privatex01  INT

11Description: How often to do 1D IOASCII output, overrides IO::outxevery

11 Range 10 x01 Default: -1
111:*  10Every so many iterations
110:  10Disable 1D IOASCII output
11-1:  10Default to IO::outxevery

11out1dxstyle  10Scope: privatex01  KEYWORD

11Description: Which style for 1D lines IOASCII output

11 Range 10 x01 Default: xgraph
11xgraph  10f over x plots suitable for xgraph
11gnuplot f(x)  10f over x plots suitable for gnuplot
11gnuplot f(t,x)  10f over t,x plots suitable for gnuplot

11

11out1dxvars  10Scope: privatex01  STRING

11Description: Variables to output in 1D IOASCII file format

11 Range 10 x01 Default: (none)
11.+  10Space-separated list of fully qualified variable/group names
11ˆ$  10An empty string to output nothing

11

11out1dxx  10Scope: privatex01  BOOLEAN

11Description: Do 1D IOASCII output in the x-direction

11                   10  x01  Default: yes

11

11out1dxxlinexy  10Scope: privatex01  REAL

11Description: y-coord for 1D lines in x-direction

11 Range 10 x01 Default: -424242
11*:*  10A value between [ymin, ymax]
11-424242:  10Default to IO::outxxlinexy

11

11out1dxxlinexyi  10Scope: privatex01  INT

11Description: y-index (from 0) for 1D lines in x-direction

11 Range 10 x01 Default: -1
110:*  10An index between [0, ny)
11-1:  10Choose the default from IO::outxxlinexyi

11

11out1dxxlinexz  10Scope: privatex01  REAL

11Description: z-coord for 1D lines in x-direction

11 Range 10 x01 Default: -424242
11*:*  10A value between [zmin, zmax]
11-424242:  10Default to IO::outxxlinexz

11

11out1dxxlinexzi  10Scope: privatex01  INT

11Description: z-index (from 0) for 1D lines in x-direction

11 Range 10 x01 Default: -1
110:*  10An index between [0, nz)
11-1:  10Choose the default from IO::outxxlinexzi

11

11out1dxy  10Scope: privatex01  BOOLEAN

11Description: Do 1D IOASCII output in the y-direction

11                   10  x01  Default: yes

11

11out1dxylinexx  10Scope: privatex01  REAL

11Description: x-coord for 1D lines in y-direction

11 Range 10 x01 Default: -424242
11*:*  10A value between [xmin, xmax]
11-424242:  10Default to IO::outxylinexx

11out1dxylinexxi  10Scope: privatex01  INT

11Description: x-index (from 0) for 1D lines in y-direction

11 Range 10 x01 Default: -1
110:*  10An index between [0, nx)
11-1:  10Choose the default from IO::outxylinexxi

11

11out1dxylinexz  10Scope: privatex01  REAL

11Description: z-coord for 1D lines in y-direction

11 Range 10 x01 Default: -424242
11*:*  10A value between [zmin, zmax]
11-424242:  10Default to IO::outxylinexz

11

11out1dxylinexzi  10Scope: privatex01  INT

11Description: z-index (from 0) for 1D lines in y-direction

11 Range 10 x01 Default: -1
110:*  10An index between [0, nz)
11-1:  10Choose the default from IO::outxylinexzi

11

11out1dxz  10Scope: privatex01  BOOLEAN

11Description: Do 1D IOASCII output in the z-direction

11                   10  x01  Default: yes

11

11out1dxzlinexx  10Scope: privatex01  REAL

11Description: x-coord for 1D lines in z-direction

11 Range 10 x01 Default: -424242
11*:*  10A value between [xmin, xmax]
11-424242:  10Default to IO::outxzlinexx

11

11out1dxzlinexxi  10Scope: privatex01  INT

11Description: x-index (from 0) for 1D lines in z-direction

11 Range 10 x01 Default: -1
110:*  10An index between [0, nx)
11-1:  10Choose the default from IO::outxzlinexxi

11

11out1dxzlinexy  10Scope: privatex01  REAL

11Description: y-coord for 1D lines in z-direction

11 Range 10 x01 Default: -424242
11*:*  10A value between [ymin, ymax]
11-424242:  10Default to IO::outxzlinexy

11

11out1dxzlinexyi  10Scope: privatex01  INT

11Description: y-index (from 0) for 1D lines in z-direction

11 Range 10 x01 Default: -1
110:*  10An index between [0, ny)
11-1:  10Choose the default from IO::outxzlinexyi

11

11out2dxdir  10Scope: privatex01  STRING

11Description: Output directory for 2D IOASCII files, overrides outxdir

11 Range 10 x01 Default: (none)
11.+  10A valid directory name
11ˆ$  10An empty string to choose the default from IO::outxdir

11out2dxevery  10Scope: privatex01  INT

11Description: How often to do 2D IOASCII output, overrides IO::outxevery

11 Range 10 x01 Default: -1
111:*  10Every so many iterations
110:  10Disable 2D IOASCII output
11-1:  10Default to IO::outxevery

11

11out2dxstyle  10Scope: privatex01  KEYWORD

11Description: Which style for 2D slices IOASCII output

11 Range 10 x01 Default: gnuplot f(x,y)
11gnuplot f(x,y)  10f over x,y plots suitable for gnuplot
11gnuplot f(t,x,y)  10f over t,x,y plots suitable for gnuplot

11

11out2dxvars  10Scope: privatex01  STRING

11Description: Variables to output in 2D IOASCII file format

11 Range 10 x01 Default: (none)
11.+  10Space-separated list of fully qualified variable/group names
11ˆ$  10An empty string to output nothing

11

11out2dxxyplanexz  10Scope: privatex01  REAL

11Description: z-coord for 2D planes in xy

11 Range 10 x01 Default: -424242
11*:*  10A value between [zmin, zmax]
11-424242:  10Default to IO::outxxyplanexz

11

11out2dxxyplanexzi  10Scope: privatex01  INT

11Description: z-index (from 0) for 2D planes in xy

11 Range 10 x01 Default: -1
110:*  10An index between [0, nz)
11-1:  10Choose the default from IO::outxxyplanexzi

11

11out2dxxzplanexy  10Scope: privatex01  REAL

11Description: y-coord for 2D planes in xz

11 Range 10 x01 Default: -424242
11*:*  10A value between [ymin, ymax]
11-424242:  10Default to IO::outxxzplanexy

11

11out2dxxzplanexyi  10Scope: privatex01  INT

11Description: y-index (from 0) for 2D planes in xz

11 Range 10 x01 Default: -1
110:*  10An index between [0, ny)
11-1:  10Choose the default from IO::outxxzplanexyi

11

11out2dxyzplanexx  10Scope: privatex01  REAL

11Description: x-coord for 2D planes in yz

11 Range 10 x01 Default: -424242
11*:*  10A value between [xmin, xmax]
11-424242:  10Default to IO::outxyzplanexx

11

11out2dxyzplanexxi  10Scope: privatex01  INT

11Description: x-index (from 0) for 2D planes in yz

11 Range 10 x01 Default: -1
110:*  10An index between [0, nx)
11-1:  10Choose the default from IO::outxyzplanexxi

11out3dxdir  10Scope: privatex01  STRING

11Description: Output directory for 3D IOASCII files, overrides IO::outxdir

11 Range 10 x01 Default: (none)
11.+  10A valid directory name
11ˆ$  10An empty string to choose the default from IO::outxdir

11

11out3dxevery  10Scope: privatex01  INT

11Description: How often to do 3D IOASCII output, overrides IO::outxevery

11 Range 10 x01 Default: -1
111:*  10Every so many iterations
110:  10Disable 3D IOASCII output
11-1:  10Default to IO::outxevery

11

11out3dxstyle  10Scope: privatex01  KEYWORD

11Description: Which style for 3D volume IOASCII output

11 Range 10 x01 Default: gnuplot f(x,y,z)
11gnuplot f(x,y,z)  10f over x,y,z plots suitable for gnuplot
11gnuplot f(t,x,y,z)  10f over t,x,y,z plots suitable for gnuplot

11

11out3dxvars  10Scope: privatex01  STRING

11Description: Variables to output in 3D IOASCII file format

11 Range 10 x01 Default: (none)
11.+  10Space-separated list of fully qualified variable/group names
11ˆ$  10An empty string to output nothing

11

11outxformat  10Scope: privatex01  STRING

11Description: Which format for IOASCII floating-point number output

11 Range 10 x01 Default: .13f
11see [1] below  10output with given precision in exponential / floating point notation

11

\^({\textbackslash}.[1]?[0-9])?[EGefg]\$

11newxfilenamexscheme 10Scope: shared from IOx01BOOLEAN

11

11outxdir  10Scope: shared from IOx01STRING

11

11outxdownsamplexx 10Scope: shared from IOx01INT

11

11outxdownsamplexy 10Scope: shared from IOx01INT

11

11outxdownsamplexz 10Scope: shared from IOx01INT

11

11outxevery  10Scope: shared from IOx01INT

11

11outxfileinfo  10Scope: shared from IOx01KEYWORD

11

11outxxlinexy  10Scope: shared from IOx01REAL

11

11outxxlinexyi  10Scope: shared from IOx01INT

11

11outxxlinexz  10Scope: shared from IOx01REAL

11

11outxxlinexzi  10Scope: shared from IOx01INT

11

11outxxyplanexz  10Scope: shared from IOx01REAL

11

11outxxyplanexzi  10Scope: shared from IOx01INT

11

11outxxzplanexy  10Scope: shared from IOx01REAL

11

11outxxzplanexyi  10Scope: shared from IOx01INT

11outxylinexx  10Scope: shared from IOx01REAL

11

11outxylinexxi  10Scope: shared from IOx01INT

11

11outxylinexz  10Scope: shared from IOx01REAL

11

11outxylinexzi  10Scope: shared from IOx01INT

11

11outxyzplanexx  10Scope: shared from IOx01REAL

11

11outxyzplanexxi  10Scope: shared from IOx01INT

11

11outxzlinexx  10Scope: shared from IOx01REAL

11

11outxzlinexxi  10Scope: shared from IOx01INT

11

11outxzlinexy  10Scope: shared from IOx01REAL

11

11outxzlinexyi  10Scope: shared from IOx01INT

11

11strictxioxparameterxcheck 10Scope: shared from IOx01BOOLEAN

11

11verbose  10Scope: shared from IOx01KEYWORD

11

6 Interfaces General Implements: ioascii

7 Schedule This section lists all the variables which are assigned storage by thorn CactusBase/IOASCII. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxSTARTUP x ioasciixstartup x startup routine x

11x11After: 11ioutilxstartup
11 x 11Language:11c
11 x 11Type: 11function

11

CCTKxBASEGRID x ioasciixchoose1d x choose 1d output lines x

11x11After: 11spatialcoordinates
11 x 11Language:11c
11 x 11Type: 11function

11

CCTKxBASEGRID x ioasciixchoose2d x choose 2d output planes x

11x11After: 11spatialcoordinates
11 x 11Language:11c
11 x 11Type: 11function